# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

### v2

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs [32, 19, 3, 10]. These models leverage cutting-edge advancements in Large Language Models (LLM) [36] and multimodal learning [25, 14, 17], enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o [10] and Claude-3.5-Sonnet [2], have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek-V3 and DeepSeek-R1 [16, 7] are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
@@
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers [29, 37]. As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses. Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms. These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models [9, 20].
@@
Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME [6], MMMU [33], MMMUPro [34], MMBench [21], PoPE [15], ChartQA [23], AI2Diagram [12] and MathVista [22]). The results highlight the substantial accuracy drop across all models when negation is introduced. The access dates for GPT-4o and Claude-3.5-Sonnet are from 3-March-2025 to 7-March-2025, while results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025. More detailed results are available in Table 1.
@@
To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across eight multimodal benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU [33] and MMBench [21], chart dataset such as ChartQA [23], to math dataset MathVista [22], providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
@@
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, suggesting potential overfitting or alignment issues during training. On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. Furthermore, different categories demonstrate varying levels of vulnerability to negation, with “Social Relation" suffering the most significant performance declines, whereas “Geography" experiences comparatively smaller yet still notable drops in our GaslightingBench. These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
2 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) [32] represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP [25] enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27] have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA [13] and Qwen2-VL [31], have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language [4, 24]. In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions [11, 5]. To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets [8]. This line of research has advanced further with studies like [28], which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding [29]. Very recent work [37] addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of [29, 37], this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) [35, 26, 30, 1]. For instance, very recent work [1] highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
@@ 3.1 Experimental Setup
Figure 5: Examples from different categories in the GaslightingBench. The highlighted green option is the correct answer, while a randomly selected incorrect option is used to generate the negation argument.
@@
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27]. These models represent the most powerful multimodal AI capabilities.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B [18], Qwen2-VL-7B-Instruct [31], LLaVA-NeXT-8B [13] and Qwen2-VL-72B-Instruct [31]. These open-source models showcase the rapid progress in publicly available multimodal AI systems.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME [6], MMMU [33], MMMUPro [34], MMBench [21] and PoPE [15], Chart dataset ChartQA [23], diagram dataset AI2Diagram [12] as well as Math dataset MathVista [22].
@@ 3.2 Evaluation Pipeline
MME [6] MMMU [33] MMMUPro [34] AI2Diagram [12] MathVista [22] ChartQA [23] PoPE [15] MMBench [21]
@@
Model LLaVA-1.6-7B [18] Qwen2-VL-7B-Instruct [31] LLaVA-NeXT-8B [13] Qwen2-VL-72B-Instruct [31] Gemini-1.5-flash [27] GPT-4o [10] Claude-3.5-Sonnet [2]
@@
Model LLaVA-1.6-7B [18] LLaVA-NeXT-8B [13] Qwen2-VL-7B-Instruct [31] Gemini-1.5-flash [27]
@@ 4.1 Performance Comparison
Comparison between Larger and Smaller Models. Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.29%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%. However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.02% decline compared to Qwen2-VL-7B-Instruct’s 32.17% drop. This suggests that larger models may be more prone to overfitting or misalignment in adversarial scenarios like negation.
@@ 5 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.

### v3

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt . These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek-V3, DeepSeek-R1 liu2024deepseek ; guo2025deepseek and Qwen series bai2023qwen ; qwen2.5 are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
@@
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning . As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses. Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms sharmatowards . These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
@@
Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista ). The results highlight the substantial accuracy drop across all models when negation is introduced. The access dates for GPT-4o and Claude-3.5-Sonnet are from 3-March-2025 to 7-March-2025, while results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025. More detailed results are available in Table 1.
@@
To comprehensively understand the limitations of MLLMs in handling negation, this paper presents the first systematic study by conducting extensive evaluations across eight multimodal benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
@@
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, indicating a lack of robustness in handling negation inputs. On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. Furthermore, different categories demonstrate varying levels of vulnerability to negation, with “Social Relation" suffering the most significant performance declines, whereas “Geography" experiences comparatively smaller yet still notable drops in our GaslightingBench. These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
@@ 2.1 Experimental Setup
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 . These models represent the most powerful multimodal AI capabilities. All evaluations for proprietary models are performed via their publicly available APIs.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B . These open-source models showcase the rapid progress in publicly available multimodal AI systems. All evaluations for these models are conducted using a single NVIDIA H100 GPU.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
@@ 2.2 Evaluation Pipeline
Figure 5: Examples from different categories in the GaslightingBench. The green-highlighted option is correct, while a randomly chosen incorrect option is used to generate the negation argument.
2.3 Gaslighting Benchmark Construction
3 Result Analysis
@@
MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
@@
Model LLaVA-1.6-7B liu2024llavanext-llava1.6 Qwen2-VL-7B-Instruct Qwen2VL-7B LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-72B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5 GPT-4o hurst2024gpt Claude-3.5-Sonnet anthropic2024claude
@@
Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
@@ 3.1 Performance Comparison
Comparison between Larger and Smaller Models. Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.29%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%. However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.02% decline compared to Qwen2-VL-7B-Instruct’s 32.17% drop. This suggests that larger models may be more prone to adversarial scenarios like negation.
@@ 3.2 Discussion: Effect of Negation Argument
4 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA li2024llavanext-strong and Qwen2-VL Qwen2VL-7B , have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications. We leave for future work the exploration of advanced reasoning foundation models, such as OpenAI’s O1 and O3, which are specifically designed to improve multi-step reasoning and tool use across modalities.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development . In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert . To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding . This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can . Recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision . For instance, recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
@@ 5 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in ensuring the trustworthiness of their outputs. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.